# S3-Attention:Attention-Aligned Endogenous Retrieval for Memory-Bounded Long-Context Inference

### v1

@@ Appendix F Extended Qualitative Examples
Sample 86. Pluralist School and Atomism. Query: The Pluralist school is said to have included what creator of the theory of atomism? Description: The core of the query is to locate the creator of atomism. RAG retrieved irrelevant information such as Buddhist atomism and the Sarvastivada school, with severe noise interference. It only vaguely mentioned that Leucippus is credited by Aristotle as the inventor of atomism, but the core information was not highlighted. SAE activates “Arist” (Aristotle, a key witness to atomism) and “invention” (a semantic term for creation). Through semantic density estimation, it filters out noise from Buddhist atomism, prioritizes the key information of Leucippus, aligns with the LLM’s reasoning logic of “Pluralist school $\rightarrow$ atomism $\rightarrow$ creator”, and accurately locks in Leucippus, solving RAG’s noise interference and lack of prominent core information.

### v2

@@ Results.
Figure 13: Historical Reasoning Samples. Sample 86. Pluralist School and Atomism. Query: The Pluralist school is said to have included what creator of the theory of atomism? Description: The core of the query is to locate the creator of atomism. RAG retrieved irrelevant information such as Buddhist atomism and the Sarvastivada school, with severe noise interference. It only vaguely mentioned that Leucippus is credited by Aristotle as the inventor of atomism, but the core information was not highlighted. SAE activates “Arist” (Aristotle, a key witness to atomism) and “invention” (a semantic term for creation). Through semantic density estimation, it filters out noise from Buddhist atomism, prioritizes the key information of Leucippus, aligns with the reasoning logic of “Pluralist school $\rightarrow$ atomism $\rightarrow$ creator”, and accurately locks in Leucippus, solving RAG’s noise interference and lack of prominent core information.